Find the equation of the line perpendicular to 3x+y-2=0 that goes through the point, (0,1)

Show all work.
Answer #1

3x+4-2=0 convert the equation into slope -intercept form ly=math- pm-slope bo b-y-intercept 37+4-2=0 3x +y = 2 Detery=-3x+2 M